Determine whether ~u and ~v are orthogonal vectors, make an acute or obtuse angle: (1.1) ~u =< 1, 3, −2 >, ~v =< −5, 3, 2 >. (2) (1.2) ~u =< 1, −2, 4 >, ~v =< 5, 3, 7 >.


1
Expert's answer
2021-07-26T15:20:28-0400

(1.1)


"u\cdot v=1(-5)+3(3)-2(2)=0"

The vectors "u" and "v" are orthogonal.


(1.2)


"u\cdot v=1(5)-2(3)+4(7)=27>0"




"=>\cos\theta>0=>0\degree<\theta<90\degree"

The vectors "u" and "v" make an acute angle.